one angle of a triangle measures 15 less than twice the measured of a second angle.the third angle measure 33. find the measure of the largest angle of the triangle

Answer by checkley79(3341)

X=2Y-15
X+Y+33=180
2Y-15+Y+33=180
3Y=180-33+15
3Y=162
Y=162/3
Y=54 ANSWER FOR THE SECOND ANGLE.
X=2*54-15
X=108-15
X=93 ANSWER FOR THE FIRST/LARGEST ANGLE.
PROOF:
54+93+33=180
180=180